OpenStudy (anonymous):

Cal 2: Please help me set up this integral. A metal company has their production increasing at a rate proportional to the product of the number of units N and the time t in years

OpenStudy (amistre64):

so a rate of change is proportional to ..... well whats your best attempt?

OpenStudy (anonymous):

dp/dn = NT? which becomes integral (1/n)dp = integral t dt

OpenStudy (amistre64):

dn/dt = nt seperates as dn/n = t dt

OpenStudy (anonymous):

is that the right integral?

OpenStudy (amistre64):

integrate both sides ln(n) = 1/2 t^2 + C ; base e it n = C e^(t^2/2)

OpenStudy (amistre64):

i cant see why it wouldnt be ...

OpenStudy (anonymous):

can i give you the other half of the question? and we could try and figure it out?

OpenStudy (amistre64):

by all means ..

OpenStudy (anonymous):

400 units are presently produced. In 2 years, 1200 units are expected. What will be the production in 4 years?

OpenStudy (amistre64):

when t=0, we are presently at 400; e^0 = 1 co C = 400 n(t) = 400 e^(t^2/2) what is n(2) ?

OpenStudy (amistre64):

err, t=4 that is

OpenStudy (anonymous):

the final answer is 32 400 units

OpenStudy (amistre64):

do we know that according to an answer key?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

dn/dt is the rate of the number of items per time frame ... does that define the rate of production to you?

OpenStudy (anonymous):

not really, i don't understand, do you get the same answer as in the answer key?

OpenStudy (amistre64):

well my 400 e^(t^2/2) did not get the same results no which is why im wondering about the setup

OpenStudy (amistre64):

never really went over modeling with differentials

OpenStudy (anonymous):

this is a differential equation

OpenStudy (anonymous):

its ment to be setup as a differential equation.

OpenStudy (anonymous):

basically with the info that is given , we must make the initial formula then make it into a integral so ex the (dn/dt) = nt then put that into a integral

OpenStudy (anonymous):

once we solve for N. We must find out the missing variables ex. C and also variable T

OpenStudy (anonymous):

once we do that, we figure out the question. In this case how long will it be in 4 years.

OpenStudy (amistre64):

it would seem to me that the production defines the number of items made, and that the rate of change of production is the rate at which the number of items is changing yeah, and we have to model it with the information given. dn/dt = nt , it seperates dn/n = t dt , it integrates ln(n) = t^2/2 + C and we solve for n n = Ce^(t^2/2) the number produced over a given time frame.

OpenStudy (anonymous):

yes now we solve for the question.

OpenStudy (anonymous):

So for ex. N(0) = 400, N(2) =1200, so what's N(4)=?

OpenStudy (amistre64):

thats the concept yes ... and i assume we use the formula from part A to account for it ...

OpenStudy (anonymous):

basically we have to find C and also T

OpenStudy (anonymous):

the missing variables

OpenStudy (amistre64):

t is given, its the number of years

OpenStudy (anonymous):

true so we need to find C.

OpenStudy (anonymous):

When we let T=0, c=400

OpenStudy (amistre64):

if our model is correct 400 = C e^0 ... C = 400

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now how do we get the answer..

OpenStudy (amistre64):

if our model is accurate 400 e^(2^2/2) = 1200, but its not ... its 2956 which leads me to thing that my thought of dn/dt isnt going to play well and we need another strategy

OpenStudy (anonymous):

i guess so.

OpenStudy (anonymous):

Well actually isnt it (E^4^2/2)?

OpenStudy (anonymous):

the final answer is 32 400 units

OpenStudy (amistre64):

lets do this dp/dt = nt see where it takes us ... consider n a constant p(t) = n t^2/2 + C , dunno how good it is but at lets see if it fails or not 400 = n + C 1200 = 2n + C maybe?

OpenStudy (anonymous):

i guess

OpenStudy (amistre64):

- 400 = -n - C 1200 = 2n + C ------------- 800 = n C = 400 p(t) = 800 t^2/2 + 400 does p(4) get us the soltion?

OpenStudy (amistre64):

6800 ... sigh, well im at a loss

OpenStudy (anonymous):

it should

OpenStudy (anonymous):

LOL

OpenStudy (amistre64):

@ rational any ideas?

OpenStudy (amistre64):

tag didnt stick lol @rational , any ideas?

OpenStudy (anonymous):

idk man, these questions confuse me

OpenStudy (amistre64):

they dont inspire me much either p = Ce^(t^2/2) + K 400 = Ce^0 + K 1200 = Ce^2 + K ---------------- 800 = C(e^2-1) C = 800/(e^2-1) K = 400 - 800/(e^2-1) p(t) = 800 e^(t^2/2)/(e^2-1) + 400 - 800/(e^2-1) p(t) = 800 (e^(t^2/2)-1)/(e^2-1) + 400 p(4) = 373 532.8 not it ....

OpenStudy (anonymous):

my teacher had a ex.

OpenStudy (anonymous):

a different question obv.

OpenStudy (anonymous):

dp/dt =Kp

OpenStudy (anonymous):

which becomes integral 1/p dp = integral k dt

OpenStudy (anonymous):

then it becomes p = e^kt + c

OpenStudy (anonymous):

how come k does become k^2/2

OpenStudy (amistre64):

k is a constant with respect to t

OpenStudy (amistre64):

unless its a function of t

OpenStudy (anonymous):

so its not the same to the question i have?

OpenStudy (amistre64):

dp/dt = kp dp/p = k dt ln(p) = kt + C p = Ce^(kt)

OpenStudy (amistre64):

C = 400 1200 = 400e^(2k) ln(3) = 2k ln(3)/2 = k try it p = 400 e^(t ln(3)/2)

OpenStudy (amistre64):

p(4) = 3600 in that case ...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=exponential+function+%280%2C400%29%2C%282%2C1200%29%2C%284%2C32400%29 now we have an issue .. there is not best fit exponential function that fits those 3 points exactly